3x-4y+9 when x=-2 and y=-5
Answer by JBarnum(2146)

So what are u trying to solve? You already have what x and y equal so what is the problem? Just substitute numbers for letters and simplify you should get 23
3*-2=-6
-4*-5=20
-6+20+9
23